# FloVD: Optical Flow Meets Video Diffusion Model for Enhanced Camera-Controlled Video Synthesis

### v1

@@ Abstract
This paper presents FloVD, a novel optical-flow-based video diffusion model for camera-controllable video generation. FloVD leverages optical flow maps to represent motions of the camera and moving objects. This approach offers two key benefits. Since optical flow can be directly estimated from videos, our approach allows for the use of arbitrary training videos without ground-truth camera parameters. Moreover, as background optical flow encodes 3D correlation across different viewpoints, our method enables detailed camera control by leveraging the background motion. To synthesize natural object motion while supporting detailed camera control, our framework adopts a two-stage video synthesis pipeline consisting of optical flow generation and flow-conditioned video synthesis. Extensive experiments demonstrate the superiority of our method over previous approaches in terms of accurate camera control and natural object motion synthesis.
@@ 1 Introduction
Figure 1: Synthesized video frames with ’zoom-out’ camera motion. X-t slice reveals pixel value changes along the red line. Our method shows natural object motion and accurate camera control, while CameraCtrl [11] produces a foreground object without motions, and MotionCtrl [33] produces artifacts.
@@
To enable realistic video synthesis with natural object motions and accurate camera control, our key idea is using optical flow as conditional input to a video diffusion model. This approach provides two key benefits. First, since optical flow can be directly estimated from videos, our method eliminates the need for using datasets with ground-truth camera parameters. This flexibility enables our model to utilize arbitrary training videos. Second, since background motion of optical flow encodes 3D correlations across different viewpoints, our method enables detailed camera control by leveraging the background motion. As a result, our approach facilitates natural object motion synthesis and precise camera control (Fig. 1).
Based on the key idea, this paper presents FloVD, a novel camera-controllable video generation framework that leverages optical flow. Given a single image and camera parameters, FloVD synthesizes future frames with a desired camera trajectory. To this end, our framework employs a two-stage video synthesis pipeline: optical flow generation and flow-conditioned video synthesis. We generate optical flow maps representing motions of the camera and moving objects from the input image and camera parameters. This flow maps are then fed into a flow-conditioned video synthesis model to generate the final output video.
To synthesize natural object motion while supporting detailed camera control, FloVD divides the optical flow generation stage into two sub-problems: camera flow generation and object flow generation. First, we convert input camera parameters into optical flow of the background motion by using 3D structure estimated from the input image. Next, an object motion synthesis model is introduced to generate the optical flow for object motions based on the input image. By combining both the background and object motion flows, we obtain the final optical flow maps.
@@ 2 Related Work
2.1 Camera-Controllable Video Synthesis
Following the tremendous success of video diffusion models, numerous efforts have been made to integrate camera controllability into the video generation process. MovieGen [22] uses text descriptions that describe camera motions to control camera motion. MCDiff [6], DragNUWA [40], and MotionI2V [25] enable camera control through user-provided strokes, manipulating background motion to adjust camera movement. AnimateDiff [9] and Direct-a-video [38] enable camera control by training models on an augmented video datasets that contain simple camera movements, such as translation. While these methods offer basic camera control with high-level instructions such as zoom and pan, they lack the detailed control capability for user-defined specific camera motions.
@@ 2.1 Camera-Controllable Video Synthesis
While these methods support detailed camera control, they are primarily trained on restricted datasets [43] due to the requirement for camera parameters during training. This limitation degrades the generalization ability and leads to unnatural object motion synthesis. In contrast, our model can be trained on arbitrary videos by leveraging optical flow maps as input, which can be robustly estimated using recent optical flow estimation models [28].
Figure 2: Overview of FloVD. Given an image and camera parameters, our framework synthesizes a video frames following the input camera trajectory. To this end, we synthesize two sets of optical flow maps that represent camera and object motions. Then, two optical flow maps are integrated and fed into the flow-conditioned video synthesis model, enabling camera-controllable video generation.
2.2 Object Motion Synthesis
Besides camera motion control, controlling object motion during video generation has been practiced either by utilizing user-provided inputs or by learning the distribution of object motion from data. Several studies employ user-stroke input [6, 40, 25, 10] or modulation of attention maps during diffusion model’s denoising process [38] to enable desired control of object motion. Another research direction focuses on synthesizing object motion by learning the diverse motions present in video datasets [19, 15, 8, 30, 7]. Given a single image, these methods synthesize flow maps to represent natural object motion and use them to animate the input image. However, these methods are primarily aimed at object motion synthesis and do not support detailed camera motion control during video generation.
@@ 3 FloVD Framework
Fig. 2 presents an overview of FloVD. Our framework takes an image $I^{s}$ and camera parameters $\mathcal{C}=\{C_{t}\}_{t=1}^{T}$ as input where $t$ is a video frame index, and $T$ is the number of video frames. $C_{t}$ is defined as a set of extrinsic and intrinsic camera parameters. Given the input conditions, our framework synthesizes a video =$\mathcal{I}=\{I_{t}\}^{T}_{t=1}$ that starts from $I^{s}$ as the first frame and follows the input camera trajectory, where $I_{t}$ is the $t$-th video frame.
FloVD consists of two stages. First, the flow generation stage synthesizes two sets of optical-flow maps that represent camera and object motions using 3D warping and an object motion synthesis model (OMSM), respectively. We refer to these optical flow maps as camera flow maps and object flow maps. These flow maps are integrated to form a single set of optical flow maps, which we refer to as camera-object flow maps. In the subsequent stage, a flow-conditioned video synthesis model (FVSM) synthesizes a video using the input image $I^{s}$ and the camera-object flow maps. In the following, we describe each stage in detail.
@@ 3.1 Flow Generation
In the flow generation stage, we first generate camera flow maps $\mathcal{F}^{c}=\{f^{c}_{t}\}_{t=1}^{T}$ where $f^{c}_{t}$ is an optical flow map from the first frame to the $t$-th frame. To generate camera flow maps reflecting the 3D structure in the input image $I^{s}$, we estimate a depth map $d^{s}$ from $I^{s}$ using an off-the-shelf single-image 3D estimation network [37]. Using the estimated depth map, we unproject each pixel coordinate $x^{s}$ in the input image $I^{s}$ into the 3D space. Then, for each $t$, we warp the unprojected coordinates and project them back to the 2D plane using $C_{t}$ to obtain the warped coordinate $x_{t}$. Finally, we construct the camera flow map $f_{t}^{c}$ by computing displacement vectors from $x^{s}$ to $x_{t}$.
@@
In this stage, we also generate object flow maps $\mathcal{F}^{o}=\{f_{t}^{o}\}_{t=1}^{T}$ that represent object motions independent of background motions. To this end, we develop OMSM based on the latent video diffusion model [3]. Specifically, as shown in Fig. 3(a), OMSM consists of a denoising U-Net, and an encoder and decoder of the latent video diffusion model’s variational autoencoder (VAE). In OMSM, the input image $I^{s}$ is first encoded by the VAE encoder. Then, the denoising U-Net takes a concatenation of the encoded input image and a noisy latent feature volume as input, and iteratively denoises the latent feature volume to synthesize latent object motion flow maps. Finally, the VAE decoder decodes the synthesized result and produces object flow maps $\mathcal{F}^{o}$.
@@
First, we estimate a binary mask $M^{obj}$ from the input image $I^{s}$ using an off-the-shelf segmentation model [23], which indicates pixels corresponding to moving objects. We use a single binary mask $M^{obj}$ for $t$, as all flow maps are forward directional optical flow maps from the first frame to the $t$-th frame. Based on $M^{obj}$, we combine $f_{t}^{c}$ and $f_{t}^{o}$. Specifically, for each pixel $x$ specified by $M^{obj}$, we compute its displaced position $x^{\prime}$ using the object motion in $f_{t}^{o}$ as $x^{\prime}=x+f_{t,x}^{o}$, where $f_{t,x}^{o}$ is the optical flow vector in $f_{t}^{o}$ at pixel $x$. Next, we transform $x^{\prime}$ using the camera parameter $C_{t}$ and the depth map $d^{s}$, obtaining $x^{\prime}_{t}$, which represents the displaced position of $x$ at the $t$-th frame due to both camera and object motions. We then compute the flow vector $f^{\prime}_{t,x}=x^{\prime}_{t}-x$. Finally, we derive the camera-object flow map $f_{t}$ as:
@@
Our OMSM is trained to generate non-zero flow vectors only for dynamic objects. Therefore, we do not necessarily need to use the mask $M^{obj}$ in Eq. 1; instead, we can transform the entire object motion flow maps using the camera parameters. However, we found empirically that using the mask $M^{obj}$ improves video synthesis quality by removing incorrectly synthesized flow vectors in static regions of $f_{t}^{o}$.
@@ 3.2 Flow-Conditioned Video Synthesis
The flow-conditioned video synthesis stage synthesizes a video $\mathcal{I}$ using the input image $I^{s}$ and the camera-object flow maps $\mathcal{F}$ as conditions. To achieve this, our framework utilizes FVSM, which extends the latent video diffusion model [3] by incorporating an additional flow encoder inspired by the T2I-Adapter architecture [20]. Specifically, as shown in Fig. 3(b), our model consists of a flow encoder $E^{f}$, a denoising U-Net, and a VAE encoder and decoder.
@@
where $\xi_{(t,l)}$ is a flow embedding of the $t$-th frame $I_{t}$ at level $l$. Each flow embedding has the resolution of its corresponding layer’s latent feature in the denoising U-Net. The denoising U-Net takes a concatenation of the encoded input image and a noisy latent feature volume as input, iteratively denoises the latent feature volume of the video. Additionally, the denoising U-Net also takes the multi-level flow embeddings by adding each of them to the feature at each layer of the denoising U-Net. Finally, the synthesized video frames are obtained by decoding the denoised latent feature volume using the VAE decoder. More details on the network architecture can be found in the supplemental document.
@@ 4.2 Training Object Motion Synthesis
OMSM is trained in two stages. The first stage initializes the model with the parameters of a pre-trained video diffusion model, and trains the model on the full dataset. The second stage fine-tunes the model using the curated dataset without camera motions. During training, we update only the parameters of the denoising U-Net, while fixing the parameters of the VAE encoder and decoder. We train the model via denoising score matching [16].
@@
Figure 4: Object flow maps synthesized by OMSM trained on the full dataset (left), and the curated dataset (right). White indicates optical flow vectors with no motion.
@@ 4.3 Training Flow-Conditioned Video Synthesis
We initialize FVSM using the parameters of a pretrained video diffusion model [3]. Then, we train only the flow encoder while fixing the other components. Similar to OMSM, we train FVSM via denoising score matching [16]. While the optical flow maps are directly estimated from video datasets in the training time, the camera-object flow maps used in the inference time are synthesized through 3D warping and OMSM. Nevertheless, both optical flow maps contain camera and object motions in the form of flow vectors, enabling the FVSM to effectively produce natural videos with the desired camera motion.
@@ 5.1 Implementation Details
FloVD synthesizes 14 video frames at once, following Stable Video Diffusion [3]. We use a resolution of $320\times 576$ for both video frames and optical flow maps. FVSM is trained for 50K iterations with 16 video clips and their optical flow maps per training batch. OMSM is trained on the full dataset for 100K iterations and then fine-tuned using the curated dataset for 50K iterations, with 8 optical flow maps per training batch. Inspired by T2I-Adapter [20], we use a quadratic timestep sampling strategy (QTS) in training FVSM for better camera controllability (Tab. 4). For stable training and inference of FloVD, we adaptively normalize optical flow maps based on statistics computed from the training dataset, following Li et al. [19]. Refer to the supplemental document for more implementation details.
@@ 5.2 Evaluation Protocol
We evaluate the video synthesis quality in terms of (1) sample quality and (2) object motion synthesis quality. To evaluate the sample quality, we first construct a benchmark dataset using 1,500 real-world videos randomly sampled from the Pexels dataset [1], which we refer to as ‘Pexels-random.’ To evaluate model’s capability of diverse object motion synthesis, we construct three benchmark video datasets with small, medium, and large object motions, each containing 500 video clips with minimal camera motions to avoid potential bias caused by camera motion. The datasets are categorized based on the average magnitudes of the optical flow vectors of moving objects: smaller than 20 pixels (Pexels-small), between 20 and 40 pixels (Pexels-medium), and more than 40 pixels (Pexels-large). More details on the benchmark datasets can be found in the supplemental document.
To evaluate the video synthesis quality, we synthesize videos using the first frames of videos in the aformentioned benchmark datasets and compare these synthesized videos with the datasets. While our method’s video synthesis quality is minimally affected by these parameters, previous methods that synthesize video frames directly from them might be more influenced. To account for this, we utilize seven types of camera trajectories during video synthesis: translation to the left, right, up, and down, as well as zoom-in, zoom-out, and no camera motion (‘stop’). Consequently, for all models, we generate seven videos for each video included in the benchmark datasets. Finally, we evaluate the video synthesis performance of a given method through the Frechet Video Distance (FVD) [29], Frechet Image Distance (FID) [13], and Inception Score (IS) [24].

### v2

@@ Abstract
We present FloVD, a novel video diffusion model for camera-controllable video generation. FloVD leverages optical flow to represent the motions of the camera and moving objects. This approach offers two key benefits. Since optical flow can be directly estimated from videos, our approach allows for the use of arbitrary training videos without ground-truth camera parameters. Moreover, as background optical flow encodes 3D correlation across different viewpoints, our method enables detailed camera control by leveraging the background motion. To synthesize natural object motion while supporting detailed camera control, our framework adopts a two-stage video synthesis pipeline consisting of optical flow generation and flow-conditioned video synthesis. Extensive experiments demonstrate the superiority of our method over previous approaches in terms of accurate camera control and natural object motion synthesis.
Figure 1: (Left) Our method using optical flow enables video synthesis with complex camera movements (dolly zoom). (Right) Synthesized video frames with ’zoom-out’ camera motion. X-t slice reveals pixel value changes along the red line. Our method shows natural object motion and accurate camera control, while CameraCtrl [9] produces an object without motions, and MotionCtrl [33] produces artifacts.
@@ 1 Introduction
To enable natural object motion synthesis and accurate camera control, our key idea is to use optical flow as conditional input to a video diffusion model. This approach provides two key benefits. First, since optical flow can be directly estimated from videos, our method eliminates the need for using datasets with ground-truth camera parameters. This flexibility enables the utilization of arbitrary training videos. Second, since background optical flow encodes 3D correlations across different viewpoints, our method enables detailed camera control by leveraging the background motion. As a result, our approach facilitates natural object motion synthesis and precise camera control (Fig. 1).
Based on the key idea, this paper presents FloVD, a novel camera-controllable video generation framework that leverages optical flow. Given a single image and camera parameters, FloVD synthesizes future frames with a desired camera trajectory. To this end, our framework employs a two-stage video synthesis pipeline: optical flow generation and flow-conditioned video synthesis. We generate optical flow maps representing the motions of the camera and moving objects from the input image and camera parameters. These flow maps are then fed into a flow-conditioned video synthesis model to generate the final output video.
To synthesize natural object motion while supporting detailed camera control, FloVD divides the optical flow generation stage into two sub-problems: camera flow generation and object flow generation. First, we convert input camera parameters into optical flow of the background motion by using 3D structure estimated from the input image. Next, an object motion synthesis model is introduced to generate the optical flow for object motions based on the input image. We obtain the final optical flow maps by combining the background and object motion flows.
@@ 2 Related Work
Camera-controllable video synthesis.
Following the tremendous success of video diffusion models, numerous efforts have been made to integrate camera controllability into the video generation process. MovieGen [23] uses text descriptions that describe camera motions to control camera motion. MCDiff [6], DragNUWA [39], and MotionI2V [26] enable camera control through user-provided strokes, manipulating background motion to adjust camera movement. AnimateDiff [8] and Direct-a-video [37] enable camera control by training models on augmented video datasets that contain simple camera movements, such as translation. While these methods offer basic camera control with high-level instructions such as zoom and pan, they lack the detailed control capability for user-defined specific camera motions.
@@
While these methods support detailed camera control, they are primarily trained on restricted datasets [43] due to the requirement for camera parameters during training. This limitation degrades the generalization capability and leads to unnatural object motion synthesis. In contrast, our model can be trained on arbitrary videos by leveraging optical flow maps as input, which can be robustly estimated using recent optical flow estimation models [29].
Figure 2: Overview of FloVD. Given an image and camera parameters, our framework synthesizes video frames following the input camera trajectory. To this end, we synthesize two sets of optical flow maps that represent camera and object motions. Then, two optical flow maps are integrated and fed into the flow-conditioned video synthesis model, enabling camera-controllable video generation.
Flow-based two-stage video synthesis.
Recently, several studies have introduced optical-flow-based two-stage pipelines for video synthesis [17, 13, 7, 19, 41, 21, 18]. These approaches, similar to ours, utilize two distinct models: one to generate optical flow maps and another to produce video frames based on the generated optical flow. However, these approaches aim to improve video synthesis quality in terms of temporal coherence, and do not address camera-controlled video synthesis. Furthermore, they do not distinguish between camera and object motions, thus extending them to incorporate camera control is not straightforward.
@@ 3 FloVD Framework
Fig. 2 presents an overview of FloVD. Our framework takes an image $I_{s}$ and camera parameters $\mathcal{C}=\{C_{t}\}_{t=1}^{T}$ as input where $t$ is a video frame index, and $T$ is the number of video frames. $C_{t}$ is defined as a set of extrinsic and intrinsic camera parameters. Given the input conditions, our framework synthesizes a video =$\mathcal{I}=\{I_{t}\}^{T}_{t=1}$ that starts from $I_{s}$ as the first frame and follows the input camera trajectory, where $I_{t}$ is the $t$-th video frame.
FloVD consists of two stages. First, the flow generation stage synthesizes two sets of optical-flow maps that represent camera and object motions using 3D warping and an object motion synthesis model (OMSM), respectively. We refer to these optical flow maps as camera flow maps and object flow maps. These flow maps are integrated to form a single set of optical flow maps, which we refer to as camera-object flow maps. In the subsequent stage, a flow-conditioned video synthesis model (FVSM) synthesizes a video using the input image $I_{s}$ and the camera-object flow maps. In the following, we describe each stage in detail.
@@ 3.1 Flow Generation
In the flow generation stage, we first generate camera flow maps $\mathcal{F}^{c}=\{f^{c}_{t}\}_{t=1}^{T}$ where $f^{c}_{t}$ is an optical flow map from the first frame to the $t$-th frame. To generate camera flow maps reflecting the 3D structure in the input image $I_{s}$, we estimate a depth map $d_{s}$ from $I_{s}$ using an off-the-shelf single-image 3D estimation network [36]. Using the estimated depth map, we unproject each pixel coordinate $x_{s}$ in the input image $I_{s}$ into the 3D space. Then, for each $t$, we warp the unprojected coordinates and project them back to the 2D plane using $C_{t}$ to obtain the warped coordinate $x_{t}$. Finally, we construct the camera flow map $f_{t}^{c}$ by computing displacement vectors from $x_{s}$ to $x_{t}$.
@@
In this stage, we also generate object flow maps $\mathcal{F}^{o}=\{f_{t}^{o}\}_{t=1}^{T}$ that represent object motions independent of background motions. To this end, we develop OMSM based on the latent video diffusion model [3]. Specifically, as shown in Fig. 3(a), OMSM consists of a denoising U-Net, and an encoder and decoder of the latent video diffusion model’s variational autoencoder (VAE). In OMSM, the input image $I_{s}$ is first encoded by the VAE encoder. Then, the denoising U-Net takes a concatenation of the encoded input image and a noisy latent feature volume as input, and iteratively denoises the latent feature volume to synthesize latent object motion flow maps. Finally, the VAE decoder decodes the synthesized result and produces object flow maps $\mathcal{F}^{o}$.
@@
First, we estimate a binary mask $M^{obj}$ from the input image $I_{s}$ using an off-the-shelf segmentation model [24], which indicates pixels corresponding to moving objects. We use a single binary mask $M^{obj}$ for $t$, as all flow maps are forward-directional optical flow maps from the first frame to the $t$-th frame. Based on $M^{obj}$, we combine $f_{t}^{c}$ and $f_{t}^{o}$. Specifically, for each pixel $x$ specified by $M^{obj}$, we compute its displaced position $x^{\prime}$ using the object motion in $f_{t}^{o}$ as $x^{\prime}=x+f_{t,x}^{o}$, where $f_{t,x}^{o}$ is the optical flow vector in $f_{t}^{o}$ at pixel $x$. Next, we transform $x^{\prime}$ using the camera parameter $C_{t}$ and the depth map $d_{s}$, obtaining $x^{\prime}_{t}$, which represents the displaced position of $x$ at the $t$-th frame due to both camera and object motions. We then compute the flow vector $f^{\prime}_{t,x}=x^{\prime}_{t}-x$. Finally, we derive the camera-object flow map $f_{t}$ as:
@@
Our OMSM is trained to generate non-zero flow vectors only for dynamic objects. Therefore, we do not necessarily need to use the mask $M^{obj}$ in Eq. 1; instead, we can transform the entire object motion flow maps using the camera parameters. However, we empirically found that using the mask $M^{obj}$ improves video synthesis quality by removing incorrectly synthesized flow vectors in static regions of $f_{t}^{o}$.
@@ 3.2 Flow-Conditioned Video Synthesis
The flow-conditioned video synthesis stage synthesizes a video $\mathcal{I}$ using the input image $I_{s}$ and the camera-object flow maps $\mathcal{F}$ as conditions. To achieve this, our framework utilizes FVSM, which extends the latent video diffusion model [3] by incorporating an additional flow encoder inspired by the T2I-Adapter architecture [20]. Specifically, as shown in Fig. 3(b), our model consists of a flow encoder $E^{f}$, a denoising U-Net, and a VAE encoder and decoder.
@@
where $\xi_{(t,l)}$ is a flow embedding of the $t$-th frame $I_{t}$ at level $l$. Each flow embedding has the resolution of its corresponding layer’s latent feature in the denoising U-Net. The denoising U-Net takes a concatenation of the encoded input image and a noisy latent feature volume as input, and iteratively denoises the latent feature volume of the video. Additionally, the denoising U-Net also takes the multi-level flow embeddings by adding each of them to the feature at each layer of the denoising U-Net. Finally, the synthesized video frames are obtained by decoding the denoised latent feature volume using the VAE decoder. More details on the network architecture can be found in the supplemental document.
@@ 4.2 Training Object Motion Synthesis
OMSM is trained in two stages. The first stage initializes the model with the parameters of a pre-trained video diffusion model, and trains the model on the full dataset. The second stage fine-tunes the model using the curated dataset without camera motions. During training, we only update the parameters of the denoising U-Net, while fixing the parameters of the VAE encoder and decoder. We train the model via denoising score matching [14].
@@
Figure 4: Object flow maps synthesized by OMSM, which is trained on the full dataset (left) and the curated dataset (right), respectively. White indicates optical flow vectors with no motion.
@@ 4.3 Training Flow-Conditioned Video Synthesis
We initialize FVSM using the parameters of a pretrained video diffusion model [3]. Then, we train only the flow encoder while fixing the other components. Similar to OMSM, we train FVSM via denoising score matching [14]. While the optical flow maps are directly estimated from video datasets in the training time, the camera-object flow maps used in the inference time are synthesized through 3D warping and OMSM. Nevertheless, both optical flow maps contain camera and object motions in the form of flow vectors, enabling the FVSM to produce natural videos with the desired camera motion effectively.
@@ 5.1 Implementation Details
FloVD synthesizes 14 video frames at once, following Stable Video Diffusion [3]. We use a resolution of $320\times 576$ for both video frames and optical flow maps. FVSM is trained for 50K iterations with 16 video clips and their optical flow maps per training batch. OMSM is trained on the full dataset for 100K iterations and then fine-tuned using the curated dataset for 50K iterations, with 8 optical flow maps per training batch. Inspired by T2I-Adapter [20], we use a quadratic timestep sampling strategy (QTS) in training FVSM for better camera controllability (Tab. S2 in the supplemental document). For stable training and inference of FloVD, we adaptively normalize optical flow maps based on statistics computed from the training dataset, following Li et al. [17]. Refer to the supplemental document for more implementation details.
@@ 5.2 Evaluation Protocol
We evaluate the camera controllability following previous methods [9, 42]. For an input image and camera parameters, we first synthesize a video. We then estimate camera parameters from the synthesized video using GLOMAP [22], and compare them against the input camera parameters to evaluate how faithfully the synthesized video follows the input parameters. For the evaluation, we sampled 1,000 video clips and their associated camera parameters from the test set of RealEstate10K [43].
@@
We evaluate the video synthesis quality in terms of (1) sample quality and (2) object motion synthesis quality. For sample quality, we first construct a benchmark dataset using 1,500 videos randomly sampled from the Pexels dataset [1] (Pexels-random). For the model’s capability of diverse object motion synthesis, we construct three benchmark video datasets with small, medium, and large object motions, each containing 500 video clips with minimal camera motions to avoid potential bias caused by camera motion. The datasets are categorized based on the average magnitudes of the optical flow vectors of moving objects: smaller than 20 pixels (Pexels-small), between 20 and 40 pixels (Pexels-medium), and more than 40 pixels (Pexels-large). More details on the benchmark datasets can be found in the supplemental document.
To evaluate the video synthesis quality, we synthesize videos using the first frames of videos in the aforementioned benchmark datasets and compare these synthesized videos with the datasets. While our method’s video synthesis quality is minimally affected by these parameters, previous methods that synthesize video frames directly from them might be more influenced. To account for this, we utilize seven types of camera trajectories during video synthesis: translation to the left, right, up, and down, as well as zoom-in, zoom-out, and no camera motion (‘stop’). Consequently, for all models, we generate seven videos for each video included in the benchmark datasets. Finally, we evaluate the video synthesis performance of a given method through the Frechet Video Distance (FVD) [30], Frechet Image Distance (FID) [11], and Inception Score (IS) [25].